OpenStudy (anonymous):

Find F inverse prime at a=4 f(x)=2x^3+3x^2+7x+4

OpenStudy (anonymous):

\[(f^{-1})\prime(4)\]

OpenStudy (anonymous):

\[f(x)=2x ^{3}+3x ^{2}+7x+4\]

OpenStudy (anonymous):

let: \(\large y=2x^3+3x^2+7x+4 \rightarrow x=2y^3+3y^2+7y+4 \) can you do implicit differentiation to get y' ???

OpenStudy (anonymous):

I use implicit diferentiation okay I will try

OpenStudy (anonymous):

But I get confuesed on how to get y by itself

OpenStudy (anonymous):

you mean y' by itself... ???

OpenStudy (anonymous):

yes that is what I mean

OpenStudy (anonymous):

ok... let's work this out...

OpenStudy (anonymous):

I can do the implicit part

OpenStudy (anonymous):

\(\large x=2y^3+3y^2+7y+4 \) \(\large 1=6y^2y'+6yy'+7y' \) factor out a y' on the right hand side: \(\large 1=(6y^2+6y+7)y' \)

OpenStudy (anonymous):

oooh

OpenStudy (anonymous):

so \(\large y'=\frac{1}{6y^2+6y+7} \)

OpenStudy (anonymous):

now this is the derivative of the INVERSE of f....

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

now there is some formula right?

OpenStudy (anonymous):

so can you use this to answer your question?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

I am figuring it out

OpenStudy (anonymous):

i'm sorry.. my internet is being stupid....

OpenStudy (anonymous):

since you know \(\large f^{-1}(4)=0 \), just plug in y=0 into the derivative: \(\large y'=\frac{1}{6y^2+6y+7} \)....

OpenStudy (noelgreco):

The slope of the line tangent to the inverse function at point a, is simply the reciprocal of the slope of the tangent to the function at a.